What's the present value of $1,325 discounted back 5 years if the appropriate interest rate is 6%, compounded monthly? Select the correct answer. a. $982.32 b. $994.62 c. $990.52 d. $986.42 e. $978.22
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

1325=P*(1+0.06/12)^(12*5)

P=1325/(1+0.06/12)^(12*5)

=$1325*0.741372196

=$982.32(Approx).